# Behavior-Inspired Neural Networks for Relational Inference

### v2

@@ A Datasets
In this section we provide data simulation and post processing steps for the dataset used in Section 5. Trajectories are simulated forward in time with a numerical integrator for a certain timestep and length, all given in Table 3. To reduce the dimensionality of the training data while preserving the dynamics of the system, the simulated trajectories are coarsen according the the frequency given in Table 4 to generate the final training data. We note that since $\Delta t$ is passed into RROD as a fixed hyperparameter for training, the timestep of $1.0$ is arbitrarily chosen for TrajNet++.
A.1 Pendulum
We generate a synthetic dataset of pendulum motion where pendulum dynamics are defined by the second order equation,
$\ddot{\theta}=-\frac{g}{\ell}\sin\theta,$ (23)
with $\ell=1.0$, $g=9.81$, and initial conditions sampled from the uniform distribution, $\theta_{0},\dot{\theta}_{0}\sim\mathcal{U}\left(-0.5\pi,0.5\pi\right)$. The dataset consists of 50000 training trajectories, and 12500 validation and testing trajectories, simulated subject to parameters given in Table 3.
A.2 Double pendulum
We generate a synthetic dataset of double pendulum motion where double pendulum dynamics are defined by the second order equations in [57]. The pendulum arms are defined $\ell_{1}=\ell_{2}=1.0$, pendulum bobs are defined $m_{1}=m_{2}=1.0$, gravity is defined $g=9.81$ and the initial conditions are sampled from the normal distribution, i.e., $\theta_{1,0},\dot{\theta}_{1,0},\theta_{2,0},\dot{\theta}_{2,0}\sim\mathcal{N}\left(0,0.5\right)$. The dataset consists of 50000 training trajectories, and 12500 validation and testing trajectories, simulated subject to parameters given in Table 3.
A.3 Mass-spring
We generate a synthetic dataset of mass spring motion where the mass spring dynamics are defined by the second order linear equation,
where the spring constant $k_{ij}$ is either $2.5$ or 00, and $k_{ij}=k_{ji}$. The initial conditions are sampled from the normal distribution, $\mathbf{r}_{i,0},\dot{\mathbf{r}}_{i,0}\sim\mathcal{N}\left(0,0.3\right)$. The dataset consists of 50000 training trajectories, and 12500 validation and testing trajectories, simulated subject to parameters given in Table 3.
A.4 Kuramoto oscillator
We generate a synthetic dataset of Kuramoto oscillator motion where the oscillator dynamics are defined by the first-order nonlinear equation,
where the coupling constant $k_{ij}$ is either $2.5$ or 00 and $k_{ij}=k_{ji}$. The initial conditions are sampled from the uniform distribution, $\phi_{i,t=0}\sim\mathcal{U}\left(0,2\pi\right)$. The dataset consists of 50000 training trajectories, and 12500 validation and testing trajectories, simulated subject to parameters given in Table 3.
B Training Details
In this section we provide training details on the experiments presented.
All models were trained on Dell Precision 7920 work stations. Each work station contains an Intel Xeon Gold 5220R 24 core CPU, two Nvidia A6000 GPUs, and 256GB of RAM. At training time only 1-GPU is used and the CPU and RAM is shared between two workloads. The hyperparameters used for training is given in Table 4. Furthermore the required memory for training both BINN and baseline models are given in Table 2. The reported values are taken from NVIDIA-smi [58] at the end of the first training epoch, which reports memory usage in terms of mebibyte (MiB) [59]. 1 mebibyte corresponds to 1.04858 megabyte.
Table 2: Memory requirement for RROD, NRI, and dNRI for experiments reported.
Network Mass-Spring Kuramoto TrajNet++
BINN* $496$ MiB $484$ MiB $406$ MiB
NRI $1670$ MiB $1658$ MiB $916$ MiB
dNRI $5008$ MiB $4996$ MiB $2180$ MiB
C Additional Experimental Results
Figure 9: Latent preferences for the double pendulum dataset.
Figure 10: Latent preferences for the double pendulum dataset.
C.1 Separability of Mutually Exclusive Categories
Consider nonlinear opinion dynamics given in Equation (3.1). Given the assumption that $z_{i1}=-cz_{i2}$, we can reduce terms such that belief term
$\sum_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{\mathcal{N}_{\mathrm{o}}}a^{\mathrm{o}}_{jl}z_{il}=a_{21}z_{i2}=-ca_{12}z_{i1}.$ (26)
The communication-belief term
$\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}\sum_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{\mathcal{N}_{\mathrm{o}}}a^{\mathrm{a}}_{ik}a^{\mathrm{o}}_{jl}z_{kl}$ $=\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}a^{\mathrm{a}}_{ik}a^{\mathrm{o}}_{21}z_{k2},$ (27)
$=\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}-a^{\mathrm{a}}_{ik}ca^{\mathrm{o}}_{12}z_{k1}.$ (28)
And therefore Equation (3.1) can be decoupled into Equation (1), such that
$\dot{z}_{i}=-d_{i}z_{i}+\mathcal{S}\left(u_{i}\left(\tilde{\alpha}_{i}z_{i}+\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}\tilde{a}_{ik}z_{k}\right)\right)+b_{i},$
where
$\tilde{\alpha}_{i}$ $=\alpha_{i1}-ca^{o}_{12},$ (30)
$\tilde{a}_{ik}$ $=a^{a}_{i1}-ca^{a}_{i1}a^{0}_{12}.$ (31)
C.2 Double Pendulum
We include additional observation space to preference encoding in Figure 9 and 10.
Table 3: Data generation and post processing parameters for experiments.
@@ C.2 Double Pendulum
Table 4: Hyperparameter for the experiments.
@@
Optimizer Adam [60] Adam Adam Adam Adam
@@
Stop spilling…
Stop spilling…

### v3

@@ A Datasets
A.1 Mechanical systems
All simulated mechanical system datasets were generated similarly. For each system, we generate 50,000 training trajectories, and 12,500 validation and testing trajectories. We note the numerical integrator, timestep, trajectory length and system parameters used to simulate each system in Table 3. We coarsen each simulated trajectory to reduce dataset size, and training time; the coarsening frequency for each system is provided in Table 3.
Table 3: Data generation and post processing parameters.
@@ A.1 Mechanical systems
Pendulum. We generate a synthetic dataset of pendulum motion, using the second-order dynamical equation,
$\ddot{\theta}=-\frac{g}{\ell}\sin\theta,$ (24)
with $\ell=1.0$ and gravity $g=9.81$. For each trajectory, we sample initial conditions from the uniform distribution on the interval $\left[-0.5\pi,0.5\pi\right]$.
Double pendulum. We generate a synthetic dataset of double pendulum motion, using the second-order dynamical equation,
$\left(m_{1}+m_{2}\right)l_{1}\ddot{\theta}_{1}+m_{2}l_{2}\ddot{\theta}_{2}\cos\left(\theta_{2}-\theta_{1}\right)$ $=m_{2}l_{2}\dot{\theta}_{2}^{2}\sin\left(\theta_{2}-\theta_{1}\right)-\left(m_{1}+m_{2}\right)g\sin\theta_{1},$ (25)
$l_{2}\ddot{\theta}_{2}+l_{1}\ddot{\theta}_{1}\cos\left(\theta_{2}-\theta_{1}\right)$ $=-l_{1}\dot{\theta}_{1}^{2}\sin\left(\theta_{2}-\theta_{1}\right)-g\sin\theta_{2},$ (26)
with $\ell_{1}=\ell_{2}=1.0$, $m_{1}=m_{2}=1.0$, and gravity $g=9.81$. For each trajectory, we sample initial conditions from the uniform distribution on the interval $\left[-0.5\pi,0.5\pi\right]$.
Mass-spring. We generate a synthetic dataset of mass-spring motion, using the second-order linear dynamical equation,
where the spring constant $k_{ij}$ is either $2.5$ or 00, and $k_{ij}=k_{ji}$. The initial conditions are sampled from a normal distribution with $\mu=0$ and $\sigma=0.3$.
Kuramoto oscillator. We generate a synthetic dataset of Kuramoto oscillator (Kuramoto, 1984) motion where the oscillator dynamics are defined by the first-order nonlinear equation,
where the coupling constant $k_{ij}$ is either $2.5$ or 00 and $k_{ij}=k_{ji}$. The initial conditions are sampled from a normal distribution with $\mu=0$ and $\sigma=2\pi$.
A.2 Human behavior
TrajNet++ dataset. The TrajNet++ dataset Kothari et al. (2021) is divided into real and synthetic pedestrian trajectory datasets; we use the synthetic dataset in our experiments. The synthetic dataset is comprised of $43,697$ trajectories, and each trajectory has $19$ timesteps and five agents.
NBA player movement dataset. The SportsVU dataset Linou et al. (2016) consists of NBA player trajectories captured over 631 games in the 2015/16 season. The dataset consists of $358,217$ trajectories, and each trajectory has $15$ timesteps and $11$ agents ($10$ players and the ball). We train our model on player trajectories and omit the ball.
B Training Details
In this section, we provide training details for each experiment. Our source code is publicly available at https://github.com/CAB-Lab-Princeton/Behavior-Inspired-Neural-Network.
All models were trained on Dell Precision 7920 work stations. Each work station is equipped with an Intel Xeon Gold 5220R 24 core CPU, two Nvidia A6000 GPUs, and 256GB of RAM.
We trained each model on a single GPU with CPU and RAM shared between two workloads. We report memory usage in Table 4 and training hyperparameters in Table 5. Memory usage is measured at the end of the first training epoch using NVIDIA-smi.
Table 4: Computational cost. Memory usage for BINN (ours), NRI, and dNRI is reported. Our model (BINN) has the lowest memory usage for all datasets (lower is better).
Network Mass-Spring Kuramoto TrajNet++
BINN* $\mathbf{496}$ MiB $\mathbf{484}$ MiB $\mathbf{406}$ MiB
NRI $1670$ MiB $1658$ MiB $916$ MiB
dNRI $5008$ MiB $4996$ MiB $2180$ MiB
Table 5: Experiment hyperparameters.
@@ B Training Details
Optimizer Adam Adam Adam Adam Adam
@@
C Additional Experimental Results
C.1 Consistency across random seeds
The character of the preferences learned in our BINN model is consistent across random seeds. In Figure 9, we provide supplemental observation space to preference encodings for the double pendulum system presented in Section 5. The encodings on the left were learned with a random seed of 172, and the encodings on the right were learned with a random seed of 272. The double pendulum encodings in Figure 6 were learned with a random seed of 72.
C.2 Capturing real-world human behavior
Human behavior can be less predictable than simulated systems. To assess how well our model can capture the behavior of human agents, we performed additional experiments on the SportsVU dataset Linou et al. (2016). We provide observation space to preference encodings for the SportsVU dataset Linou et al. (2016) in Figure 10. The learned relationship between preferences is similar to the learned relationship for synthetic pedestrian data (see Section 5.2.2). When the relative magnitudes of $z_{i1}$ and $z_{i2}$ change, the direction of travel changes from upwards to downwards. When the relative magnitudes of $z_{i3}$ and $z_{i4}$ change, the direction of travel changes from leftwards to rightwards.
C.3 Additional baseline comparisons
We report an additional comparison against Factorised Neural Relational Inference (fNRI) (Webb et al., 2019). Similarly to our approach, fNRI models multiple types of inter-agent relationships; however, their approach uses a factorized graph representation, while ours uses the nonlinear opinion dynamics inductive bias. We report the 10-step trajectory prediction performance of our BINN+ and fNRI on the TrajNet++ dataset in Table 6, where our BINN+ model outperforms fNRI.
Figure 9: Consistency across random seeds. We show the observed, and latent representations of each double pendulum bob. (Left-right) Encodings learned with a random seed of 172; encodings learned with a random seed of 272. (Top-bottom) The observed position and velocity of the first bob; the learned preferences of the first bob; the observed position and velocity of the second bob; the learned preferences of the second bob. When preference switch, physical bob’s motion switches direction.
Figure 10: NBA player preferences. (Left) We visualize example trajectories from the dataset. The initial position of the trajectory is marked by a blue square and the point of direction change is marked by a black square. (Right) We plot agent preferences for each latent categories. Dominant $z_{i1}$ corresponds to upward motion, and dominant $z_{i2}$ corresponds to downward motion. Dominant $z_{i3}$ corresponds to rightward motion, and dominant $z_{i4}$ corresponds to leftward motion.
Table 6: Trajectory prediction. Trajectory prediction error on the TrajNet++ datasets is reported. Our BINN+ model outperforms fNRI.
Network TrajNet++
BINN+* $\boldsymbol{3.51\times 10^{-3}}$
fNRI $6.85\times 10^{-3}$
D Aspects of Bifurcation Theory and Nonlinear Opinion Dynamics
In this section, we aim provide intuition building examples for the pitchfork bifurcation introduced in Figure 2, and the notion of mutually exclusive categories introduced in Section 3.1.
Pitchfork bifurcation. Bifurcation theory can be described as the study of equations with multiple solutions. Consider, for example, the dynamical system defined by
$\dot{z}=z^{3}-uz.$ (29)
The equilibrium solutions of this dynamical equation (i.e., the values of $z$ for which $\dot{z}=0$) are shown in the bifurcation diagram of Figure 11. When the bifurcation parameter $u$ is less than or equal to the critical value $u^{*}=0$, there is only one solution for Equation 29 (i.e., $z=0$). When $u$ is greater than $u^{*}$, the number of solutions jumps to three (i.e., $z\in\{\sqrt{u},-\sqrt{u},0\}$). The jump in the number of solutions from one to three at the critical value $u^{*}$, is the basic property that characterizes a pitchfork bifurcation. As it turns out, Equation 29 is the simplest equation with this behavior (Golubitsky et al., 2012).
The stability of solutions to the dynamical system defined in Equation (29) can be determined by evaluating the derivative
$\ddot{z}=3z^{2}-u.$ (30)
When the derivative is greater than zero, the solution is stable. When it is less than zero, the solution is unstable. For the solution, $z=0$, the derivative evaluates to
$\ddot{z}=-u,$ (31)
which is stable for $u<0$, and unstable for $u>0$. The change in the stability of the solution $z=0$ is illustrated in the bifurcation diagram of Figure 11. When $u<0$, the solution $z=0$ is illustrated with a solid line and when $u>0$, the solution is illustrated with a dotted line. The stable solutions $z\in\{\sqrt{u},-\sqrt{u}\}$ are illustrated with solid lines.
$u_{i}$$u^{*}$$z_{ij}$00
Figure 11: Pitchfork bifurcation. States are represented by $z$ and $u$ is the bifurcation parameter. When $u$ is less than the critical value $u^{*}$, there is only one solution, $z=0$. When $u$ is larger than the critical value $u^{*}$, the solutions changes to three, $z\in\{\sqrt{u},-\sqrt{u},0\}$. The stable (unstable) equilibrium solutions are illustrated with solid (dotted) lines.
In applications, the bifurcation parameter often corresponds to an external force. An example of a physical system whose solutions form a pitchfork bifurcation is a trailer being towed by a truck. In this case, the location of the center of gravity of the trailer is the bifurcation parameter and determines the stability of the truck-trailer system. A video showing the stable and unstable solutions of this physical system can be found at https://www.youtube.com/watch?v=w9Dgxe584Ss. A longer discussion of this system can be found in Horvath & Takacs (2022).
Separability of mutually exclusive categories. In the nonlinear opinion dynamics framework, the presence of mutually exclusive preferences simplifies the dynamics. Consider the case of two categories (i.e., $\mathcal{N}_{\mathrm{o}}=2$). Preferences that are opposite in sign (i.e., $z_{i1}=-cz_{i2}$) and negatively correlated (i.e., $a^{\mathrm{o}}_{12},a^{\mathrm{o}}_{21}<0$) are mutually exclusive. Given these conditions, the belief term can be simplified to
$\sum_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{\mathcal{N}_{\mathrm{o}}}a^{\mathrm{o}}_{jl}z_{il}=a_{21}z_{i2}=-ca_{12}z_{i1},$ (32)
and the communication-belief term can be simplified to
$\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}\sum_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{\mathcal{N}_{\mathrm{o}}}a^{\mathrm{a}}_{ik}a^{\mathrm{o}}_{jl}z_{kl}=\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}a^{\mathrm{a}}_{ik}a^{\mathrm{o}}_{21}z_{k2}=\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}-a^{\mathrm{a}}_{ik}ca^{\mathrm{o}}_{12}z_{k1}.$ (33)
With these simplification, the dynamics of $z_{i1}$ in Equation (1) do not depend on $z_{i2}$. Defining the decoupled self-reinforcement term
$\tilde{\alpha}_{i}=\alpha_{i1}-ca^{o}_{12},$ (34)
and communication matrix
$\tilde{a}_{ik}$ $=a^{a}_{i1}-ca^{a}_{i1}a^{0}_{12},$ (35)
Equation (1) can be simplified to Equation (2), that is,
$\dot{z}_{i}=-d_{i}z_{i}+\mathcal{S}\left(u_{i}\left(\tilde{\alpha}_{i}z_{i}+\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}\tilde{a}_{ik}z_{k}\right)\right)+b_{i}.$ (36)